OpenStudy (anonymous):

If f(x) = |(x^2 − 4)(x^2 + 2)|, how many numbers in the interval [0, 1] satisfy the conclusion of the Mean Value Theorem? a: None b: 1 c: 2 d: 3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[f'(x) = \frac{ f(b)-f(a) }{ b-a }\] MVT Your interval is [0,1] \[f(x) = |(x^2-4)(x^2+2)| = |x^4-2x^2-8| = -x^4+2x^2+8\] now you can just find the derivative using the MVT.

OpenStudy (anonymous):

i got: None! a :)